y work Tech Solutions is a consulting firm that uses a job-order costing system. Its direct materials consist of hardware and software that it purchases and installs on behalf of its clients. The firms direct labor includes salaries of consultants that work at the clients job site, and its overhead consists of costs such as depreciation, utilities, and insurance related to the office headquarters as well as the office supplies that are consumed serving clients. Tech Solutions computes its predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 80,000 direct labor-hours would be required for the periods estimated level of client service. The company also estimated $680,000 of fixed overhead cost for the coming period and variable overhead of $0.50 per direct labor-hour. The firms actual overhead cost for the year was $692,000 and its actual total direct labor was 83,000 hours. Required: 1. Compute the predetermined overhead rate. 2. During the year, Tech Solutions started and completed the Xavier Company engagement. The following information was available with respect to this job: Direct materials 38,000 Direct labor cost 21,000 Direct labor hours worked 280 Compute the total job cost for the Xavier Company engagement.

Tech Solutions is a consulting firm that uses a job-order costing system. Its direct materials consist of hardware and software that it purchases and installs on behalf of its clients. The firm's direct labor includes salaries of consultants that work at the client's job site, and its overhead consists of costs such as depreciation, utilities, and insurance related to the office headquarters as well as the office supplies that are consumed serving clients.  Tech Solutions computes its predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 80,000 direct labor-hours would be required for the period's estimated level of client service. The company also estimated exist680,000 of fixed overhead cost for the coming period and variable overhead of exist0.50 per direct labor-hour. The firm's actual overhead cost for the year was exist692,000 and its actual total direct labor was 83,000 hours.  


Required:  

1. Compute the predetermined overhead rate.  

2. During the year, Tech Solutions started and completed the Xavier Company engagement. The following information was available with respect to this job:  

Compute the total job cost for the Xavier Company engagement.
Answer #1
Concepts and reason

Job cost sheet: Job cost sheet records the details about direct materials cost, direct labor cost, and manufacturing overheads. It shows the total cost of the job work and completion details.

Variable overheads costs: These costs are directly proportional to the level of activity. If there is a change in the activity level by 10% then there would be a change in the variable cost by 10% in either direction. Variable cost is directly proportional to the volume of the production. Variable cost per unit remains constant.

Fixed overheads costs: These costs do not vary with change in the level of activity. Fixed cost remains the same for the period. If there is an increase in the activity level, then the fixed cost per unit decreases and if there is a decrease in the activity level then the fixed cost per unit increases. Fixed cost per unit changes with the change in the activity level.

Pre-determined overhead rate: It is the rate which is calculated based on the activity base. This rate is used to allocate manufacturing overheads.

Fundamentals

Calculate the predetermined overhead rate as follows:

Overheadrate=Expectedannualoverheadcost÷Expectedannualoperatingactivity{\rm{Overhead}}\,{\rm{rate}}\,{\rm{ = }}\,{\rm{Expected}}\,{\rm{annual}}\,{\rm{overhead}}\,{\rm{cost}}\, \div \,{\rm{Expected annual}}\,{\rm{operating}}\,\,{\rm{activity}}

Total job cost is calculated by adding the all cost incurred for the completion of the job i.e. direct materials cost, direct labor cost, and overheads cost.

1.

Variable overhead rate is $0.50 per direct labor hour, total estimated fixed overhead is $680,000, and the estimated direct labor hour is 80,000.

Determine the predetermined overhead rate as follows: -

PredeterminedOverheadrate=VariableOverheadRate+FixedOverheadRate=$0.50+($680,000÷80,000)=$9.00\begin{array}{c}\\\Pr e\det er\min ed\,Overhead\,rate = \,Variable\,\,Overhead\,Rate + Fixed\,Overhead\,Rate\\\\ = \,\$ 0.50 + (\$ 680,000 \div 80,000)\\\\ = \$ 9.00\\\end{array}

Therefore, the predetermined overhead rate is $9 per direct labor hour.

2.

Variable overhead rate is $0.50 per direct labor hour, total actual fixed overhead is $692,000, and the estimated direct labor hour is 83,000.

Determine the total job cost as follows: -

TotalJobCost=DirectMaterials+DirectLabor+OverheadCost=$38,000+$21,000+($9×280)=$61,520\begin{array}{c}\\Total\,Job\,\,{\mathop{\rm Cos}\nolimits} t = \,Direct\,Materials\, + Direct\,Labor\, + \,Overhead\,{\mathop{\rm Cos}\nolimits} t\\\\ = \,\$ 38,000 + \$ 21,000 + (\$ 9\, \times \,280)\\\\ = \$ 61,520\\\end{array}

Therefore, the total job cost is $61,520.

Ans: Part 1

The predetermined overhead rate is $9 per direct labor hour.

Part 2

The total job cost is $61,520.